OpenStudy (anonymous):

Last year, there were 120 seniors in Limon High School. This year there are 145. What is the percent of increase? Round to the nearest percent.

OpenStudy (anonymous):

120(1+x)=145

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

I think 20.8 %

OpenStudy (anonymous):

120+120x=145 120x=145-120

OpenStudy (anonymous):

120x= 25 x=25/120

OpenStudy (anonymous):

x=.208333 multiply by 100 to get 20.8